# Synthetic Data Applications in Finance

### 29 Dec 2023 v1

@@ 1 Introduction
Synthetic data has been receiving increased attention in the research community and beyond with the wide-spread popularity of generative models such as DALL-E [${}^{+}$21] and GPT4 [Ope23] for the domains of image and text generation, respectively. The use of synthetic data in finance is still in an early phase and typically involves tabular, time-series, and text datasets [${}^{+}$20]. In this work, we will primarily focus on tabular and time-series synthetic data which have wide applicability in the retail and investment banking applications such as marketing, trading, and anti-money laundering. Also, we will briefly touch upon modalities involving images and text as can be seen in applications such as check fraud and document understanding. The latter has seen tremendous uptick in financial use-cases with the introduction of chatGPT.
@@
Data use and sharing within and outside the financial institutions is highly restrictive due to internal policies designed to protect the important relationship of trust between consumers and financial institutions and to ensure compliance with the various regulatory regimes across the globe. These are mostly concerned about the privacy and legal aspects of the customer data and they in turn lead to limits or bureaucracy for data use. Other risks include leakage of institutional knowledge which could pose a competitive risk for the bank [LWSF23, TBV23]. Certain types of synthetic data may potentially alleviate this issue and thereby speed up the adoption of AI and model development process in the firm [${}^{+}$20]. There are many criteria for evaluating the quality of synthetic data and one such notion is that of epistemic parity [${}^{+}$22] where the findings on the original dataset match those that are found on the synthetic dataset. Other applications include explanations where instead of using samples from the real datasets, we can generate private synthetic data [PSZ22].
@@ 3.3 Defences against privacy attacks
3.4 Privacy levels: synthetic data generation techniques as privacy defenses
In this section, we propose a six-level privacy defense hierarchy and discuss the privacy attacks, utility implications, and potential privacy guarantees for each level. Each level corresponds to a category of defense mechanisms. This is important when reviewing what individuals and businesses want to do with existing data. In practice, we would look to apply the corresponding privacy level to the specific use case in order to achieve the business goal, with security and speed being the two primary benefits to the business. Note that there is still not a widely accepted definition of synthetic data [${}^{+}$22].
$\xrightarrow{\hskip 85.35826pt\text{More privacy}\hskip 85.35826pt}$
Level 1 Level 2 Level 3 Level 4 Level 5 Level 6
Obscure PII
Obscure PII
Randomization
Generative
modeling
Generative
modeling
randomization
Calibrated
simulation
Uncalibrated
Simulation
$\xleftarrow{\hskip 85.35826pt\text{More fidelity}\hskip 85.35826pt}$
Table 4: Categorization of synthetic data approaches based on privacy levels. Higher levels generally provide better privacy guarantees, but are less truthful to the original data.
@@ 3.4 Privacy levels: synthetic data generation techniques as privacy defenses
Level 1: Obscure PII
Examples of mechanisms at this level include dropping, replacing, masking, or anonymizing of the PII attributes. Since this approach does not modify non-PII attributes in any way, it will not impact the utility of downstream tasks and thus there is no utility degradation. This however represents weak privacy protection as data remains vulnerable to linkage attacks (see Figure 2 for an illustrative example).
Figure 3: Privacy Level 2: Obscure PII + Randomization
Level 2: Obscure PII + Randomization
In addition to obscuring PII columns, we add deliberate noise to other attributes or carefully randomize the algorithm applied to the data (algorithm could be a query from the dataset, train a machine learning algorithm, etc.). Depending on the amount of noise and the downstream task, some degree of utility degradation is expected (see Figure 3 for an illustrative example).
Figure 4: Privacy Level 3: Synthetisized Rows
Level 3: Generative modeling
Many ML-based synthetic data generation approaches such as Gaussian copula, GAN copula, etc. fall in this category [PWV16b, ${}^{+}$14b, ${}^{+}$18]. These methods typically do not have strong privacy guarantees (see Figure 4 for an illustrative example).
Figure 5: Privacy Level 4: Synthetisized Rows + Test
Level 4: Generative modeling + Randomization
Randomizing the synthetic data generation process by adding deliberate noise to the input data (the real data), the synthetic data generation algorithm, or the output (synthetic data). Utility degradation depends on the synthetic data generation approach, the amount of injected noise, and the downstream task. This approach has the potential to achieve strong protection against MIA. In particular, synthetic data generation methods that provide DP guarantees belong to this class [${}^{+}$19, ${}^{+}$18, YJvdS19b] (see Figure 5 for an illustrative example).
@@
Level 5: Simulation-based synthetic data generation calibrated on real data
In this approach the data generation model is not trained on real data. In fact, there is no learning in this approach. This approach relies on simulations governed by rules to generate synthetic data. These rules, however, are calibrated such that the generated data follows some statistical properties of the original, real dataset. Utility degradation depends on the downstream task and the simulation framework. This approach generally represents a strong defence, except in the context of PIA attacks, since the simulator is calibrated with respect to statistical properties of the real dataset (see Figure 6 for an illustrative example).
@@
Level 6: Uncalibrated simulation-based synthetic data generation
Uncalibrated simulation has no related statistical properties to the original real dataset, meaning that the population-level statistics from the original dataset are not revealed. In general, this method yields a strong privacy guarantee given the severe utility degradation. This method of data creation requires the user to define the metadata of the original data, describing the column names and the values in those columns. Given the uncalibrated nature of this level, different data distributions to that of the original dataset are used, strengthening the defence of the synthetic data. This approach remediates one of the consequences of level 5 generation of defence against PIA attacks, given that the statistical properties of the data are uncalibrated to the real dataset (see Figure 7 for an illustrative example).

### 20 Mar 2024 v2

@@ 1 Introduction
As the name suggests, synthetic data is artificially generated rather than produced by real world events. Synthetic data is created via two primary methods, namely: 1) By transforming real data, or 2) By simulation of real processes. We refer the reader to the rich literature on synthetic data and the many mechanisms for creating it [${}^{+}$20b]. Synthetic data has been receiving increased attention in the research community and beyond with the wide-spread popularity of generative models such as DALL-E [${}^{+}$21] and GPT4 [Ope23] for the domains of image and text generation, respectively. In this work, we will primarily focus on tabular and time-series synthetic data which have wide applicability in the retail and investment banking applications such as marketing, trading, and anti-money laundering. Also, we will briefly touch upon modalities involving images and text as can be seen in applications such as check fraud and document understanding. The latter has seen tremendous uptick in financial use-cases with the introduction of chatGPT.
@@
Data use and sharing within and outside the financial institutions is highly restrictive due to internal policies designed to protect the important relationship of trust between consumers and financial institutions and to ensure compliance with the various regulatory regimes across the globe. These are mostly concerned about the privacy and legal aspects of the customer data and they in turn lead to limits or bureaucracy for data use. Other risks include leakage of institutional knowledge which could pose a competitive risk for the bank [LWSF23, TBV23]. Certain types of synthetic data may potentially alleviate this issue and thereby speed up the adoption of AI and model development process in the firm [${}^{+}$20a]. There are many criteria for evaluating the quality of synthetic data and one such notion is that of epistemic parity [${}^{+}$22] where the findings on the original dataset match those that are found on the synthetic dataset. Other applications include explanations where instead of using samples from the real datasets, we can generate private synthetic data [PSZ22].
@@ 3.3 Defences against privacy attacks
3.4 Privacy levels
Now we introduce a six-level privacy defense hierarchy and discuss the privacy attacks, utility implications, and potential privacy guarantees for each level. Each level corresponds to a group of defense mechanisms with increasingly stronger privacy protections.
These levels can provide guidance to businesses regarding the security and utility of their Synthetic Data. For instance, they might choose to allow internal sharing of Level 2 data if it arises from a non-critical source, but require Level 4 protections for more sensitive data. The relevant privacy level should be determined according to the use case to balance multiple objectives such as the business goal, security, speed of generation, and utility.
In the first 4 levels, we consider methods where the data is transformed from the original dataset to the Synthetic Data. In the figures, the original data appears on the left, and the arrows indicate how the data is transformed. We focus on tabular data in these examples, but the principles can apply to other types of data.
@@ 3.4 Privacy levels
3.4.1 Privacy Level 1: Obscure PII
Examples of mechanisms at this level include dropping, replacing, masking, or anonymizing the PII attributes. Since this approach does not modify non-PII attributes in any way, it dones not reduce the utility of downstream tasks and accordingly there is no utility degradation. This however represents weak privacy protection as data remains vulnerable to reconstruction attacks [NS07].
Figure 3: Privacy Level 2: Obscure PII + noise
3.4.2 Privacy Level 2: Obscure PII + noise
In addition to obscuring PII columns, we can deliberately add noise to other attributes to reduce the effectiveness of potential attacks. Differential privacy techniques, for instance, can provide formal guarantees against MIA.
Another approach involves randomly “swapping” data between entries. So for instance, in a demographic dataset, the ages of the included individuals might be reordered randomly in the records. This technique aims to provide plausible but randomized data by making it more difficult for an adversary to infer any information regarding any particular individual. These techniques aim to elevate privacy while preserving the utility of the data to a downstream task.
Depending on the amount of noise and the downstream task, some degree of utility degradation is expected.
Figure 4: Privacy Level 3: Generative modeling. The question mark suggests the possibility of reverse-engineering the data.
3.4.3 Privacy Level 3: Generative modeling
Note that Privacy Levels 1 and 2 involve row-by-row transcription of the original data (with obfuscation or noise as appropriate). Accordingly, such datasets cannot be larger than the original.
With Level 3 we move to generative techniques where we analyze the original data to build a model that can create new data. Example approaches include Gaussian copula, and Generative-Adversarial-Networks (GAN) [PWV16b, ${}^{+}$14b, ${}^{+}$18]. Other methods use differential privacy techniques to offer additional guarantees [${}^{+}$19, ${}^{+}$18, YJvdS19b]. In our own work, we have introduced a KD-tree-based formulation to model the data that offers additional protections as well [${}^{+}$23].
All these methods enable the creation of new data elements distinct from the original data. They offer stronger protection than in Level 1 or Level 2, but are still potentially subject to attack. The risk is increased when the relative size of the generated data to the original data is large: For example, if we generate one million samples using an original dataset of only 1,000 we would expect to see generated samples clustering around the samples in the original data.
Figure 5: Privacy Level 4: Generative modeling + testing
3.4.4 Level 4: Generative modeling + testing
For Level 4 we add explicit testing of each generated dataset to validate its resistance to specific attacks. The particular tests and the corresponding scores required to “pass” depend on the data and the application. For instance, it may be acceptable for certain properties of the data to “leak” while others should not. To operationalize this, we leverage published attack algorithms, then score the data depending on the success of the attack.
While it is hard to specify which test and which score would be necessary to achieve Level 4 privacy in all cases, the important and critical difference above Level 3, is the fact that the data is explicitly tested. The test and the scoring criteria must be determined by the individual business for the use case. Example scoring criteria measure resistance to membership inference, attribute reconstruction, and property attacks. among others [GBWT23, ${}^{+}$22, ${}^{+}$22, ${}^{+}$23, DL24].
@@ 3.4.4 Level 4: Generative modeling + testing
3.4.5 Level 5: Calibrated simulation
In this approach the generation method is not trained on real data. In fact, there is (usually) no learning in this approach. Instead, we rely on simulations governed by rules or knowledge of the process that would otherwise generate real data. These rules, however, are calibrated with reference to the real process such that the generated data follows some statistical properties of the original, real system. As an example, we might use a simulation of the stock market to generate stock price data. In our own work, we have developed calibrated simulations of equity markets that correspond to Level 5 privacy [${}^{+}$19].
Utility degradation depends on the downstream task and the simulation framework. This approach generally represents a strong defense against adversarial attacks. However, they may be exposed to Property Inference Attacks, because the simulator is calibrated with respect to statistical properties of the real system.
@@ 3.4.5 Level 5: Calibrated simulation
3.4.6 Level 6: Uncalibrated simulation
In this case we may not be aware of the statistical properties of the modeled system, or we might deliberately avoid adjusting the simulation to correspond to the properties of the original system. Even though such a simulation may not provide high fidelity data, it can still prove quite useful.
An important use is testing, in which we might use a simulation to explore all the potential values of data fields to see if they “break” our downstream processes. Additionally, we might choose to embed known examples of situations we want to be sure our systems detect (e.g., fraudlent transactions). Another use is to create what-if scenarios where we hypothesize the impact of one factor on another, to see if visualization techniques might enable us to discover those relationships in practice.
In general, this method yields a strong privacy guarantee. It remediates one of the consequences of level 5 generation of defence against PIA attacks, given that the statistical properties of the data is uncalibrated to the real dataset.